Keyboard functions for dviscr and dvivik

^=Ctrl  ~=Alt

^C or Q exit
← left
→ right
↑ up
↓ down
4 screen left
6 screen right
8 screen up
2 or SpaceBar screen down
Home go home (:-)
7 upper left
End or 1 bottom left
^← left margin
^→ right margin
+ zoom in
- zoom out
~G nm gray scale nm
~S nm scaling nm
T n transformation n
K toggle coordinate transformation

PageUp or 9 or F7 previous page
PageDn or 3 or F8 next page
^PageUp first page
^PageDn last page
G n go to page n
O go to ‘old’ page
P n go to page number n

S text search text

Esc terminate search
H set home
F finer movement
C coarser movement
R toggle ruler
^R toggle ruler scale
W toggle window/ruler movement
U units forward
^U units backward
I inverse screen
D refresh screen